# Walking the Weight Manifold: a Topological Approach to Conditioning Inspired by Neuromodulation

### v1

@@ 2.1 Weights as parameterized functions
In traditional neural networks, each weight (synapse) is described by a single value. Here, we make each weight a function of the conditioning variable. Formally, consider a smooth manifold $\mathcal{M}(s,\mathbf{P})$ parametrized by $s\in[0,1]$ and depending on parameters $\mathbf{P}\in\mathbb{R}^{p}$. The parameter $s$ represents the modulator or conditioning value, and the parameters $\mathbf{P}$ are learnable. Then, :×$\mathcal{M}:\mathbb{R}\times\mathbb{R}^{p}\to\mathbb{R}^{d}$ is map from points on the manifold, selected by $s$, to the weights of the network.
@@ 2.3 Practical Implementation: Efficient Optimization of Weight Manifolds
Conveniently, most learnable operations in neural networks (e.g. fully-connected layers, convolutions, embeddings, sub-operations of self-attention, etc.) are also linear in that they are additive and homogenous—$f(\alpha a+\beta b)=\alpha f(a)+\beta f(b)$. Nonlinear operations such as ReLU generally do not contain any learnable components. This property along with Eq. 12 allows us to process all $B$ examples in a batch while only ever holding the $n$ basis points in memory.
@@ 3.1 Generalization to unseen conditions
Here, we design a setting where data augmentation is applied the inputs to the network, and the conditioning describes the augmentation that was applied. Specifically, we rotate CIFAR-10 images passed to a moderately sized convolutional network (CNN) and the conditioning, $s$, encodes the angle of the rotation from $[0,2\pi]$. The network must classify these images despite the rotation. To test generalization, we only sample a fixed subset of the possible angles during training (e.g. 10% of $[0,2\pi]$). During evaluation, we sample the full range and report the test accuracy as a function of the sparsity of the training conditions. Fig. 2a illustrates the overall setting.
@@ A.1 Theorem 1: Optimal Manifold Perturbation
[…]
@@ B.2.1 Metric Tensor Derivation
[…]
@@ B.2.3 Definition of Cubic B-Spline Basis Functions
A key property is that $B_{i}(s)B_{j}(s)=0$ if $|i-j|>3$, meaning the metric tensor has a banded structure with bandwidth 3.
@@ B.2.4 Integration of B-Spline Products
[…]

### v2

@@ 2.1 Weights as parameterized functions
In traditional neural networks, each weight (synapse) is described by a single value. Here, we make each weight a function of the conditioning variable. Formally, consider a smooth manifold $\mathcal{M}(s,\mathbf{P})$ parametrized by $s\in[0,1]$ and depending on parameters $\mathbf{P}\in\mathbb{R}^{p}$. The parameter $s$ represents the modulator or conditioning value, and the parameters $\mathbf{P}$ are learnable. Then, $\mathcal{M}:\mathbb{R}\times\mathbb{R}^{p}\to\mathbb{R}^{d}$ is map from points on the manifold, selected by $s$, to the weights of the network.
@@ 2.3 Practical Implementation: Efficient Optimization of Weight Manifolds
[…]
@@ 3.1 Generalization to unseen conditions
Here, we design a setting where data augmentation is applied the inputs to the network, and the conditioning describes the augmentation that was applied. Specifically, we rotate CIFAR-10 images passed to a moderately sized convolutional network (CNN) and the conditioning, $s$, encodes the angle of the rotation from $0,2\pi]$. The network must classify these images despite the rotation. To test generalization, we only sample a fixed subset of the possible angles during training (e.g. 10% of $0,2\pi]$). During evaluation, we sample the full range and report the test accuracy as a function of the sparsity of the training conditions. Fig. 2a illustrates the overall setting.
@@ A.1 Theorem 1: Optimal Manifold Perturbation
Theorem 1. Given a parametrized manifold $\mathcal{M}(s,\mathbf{P})$ with loss functional $L[\mathcal{M}]=\int_{0}^{1}\ell(\mathcal{M}(s,\mathbf{P}))ds$, the optimal perturbation $\Delta\mathbf{P}$ that minimizes the first-order approximation of $L[\mathcal{M}(s,\mathbf{P}+\Delta\mathbf{P})]$ subject to a constraint on the total squared distance $\int_{0}^{1}|\mathcal{M}(s,\mathbf{P}+\Delta\mathbf{P})-\mathcal{M}(s,\mathbf{P})|^{2}\>\mathrm{d}s=C_{0}$ is given by:
@@ B.2.1 Metric Tensor Derivation
To apply our manifold optimization approach, we need to compute the integrated metric tensor $\mathbf{\mkern 3.0mu\overline{\mkern-3.0muM\mkern-3.0mu}\mkern 3.0mu}=\int_{0}^{1}\mathbf{M}(s)ds$. The Jacobian matrix is:
@@ B.2.3 Definition of Cubic B-Spline Basis Functions
A key property is that $B_{i}(s)B_{j}(s)=0$ if |ij|>3|i-j|>3, meaning the metric tensor has a banded structure with bandwidth 3.
@@ B.2.4 Integration of B-Spline Products
$\mathbf{\mkern 3.0mu\overline{\mkern-3.0muM\mkern-3.0mu}\mkern 3.0mu}_{ij}=\int_{0}^{1}B_{i}(s)B_{j}(s)ds\cdot\mathbf{I}=\begin{cases}\frac{1}{140}\cdot\mathbf{I},&\text{if }|i-j|=3\\ \frac{1}{60}\cdot\mathbf{I},&\text{if }|i-j|=2\\ \frac{11}{140}\cdot\mathbf{I},&\text{if }|i-j|=1\\ \frac{1}{20}\cdot\mathbf{I},&\text{if }|i-j|=0\end{cases}$ (38)